OpenStudy (anonymous):

If (x + yi) + (2(x + yi) + 3i) = 9, what is x + yi?

OpenStudy (anonymous):

3 − i -3 + i -3 − i 3 + i 9 − 3i

OpenStudy (aum):

On the left hand side, group the real terms together and the imaginary terms together first. What do you get?

OpenStudy (anonymous):

What should it look like

OpenStudy (aum):

(x + yi) + (2(x + yi) + 3i) = 9 x + yi + 2x + 2yi + 3i = 9 x + 2x + (yi + 2yi + 3i) = 9 3x + (3yi + 3i) = 9 3x + (y + 3)*3i = 9 3x = 9 and y + 3 = 0 x = 3 and y = -3 x + yi = 3 - 3i

OpenStudy (aum):

None of the answers match. Can you post a link to the original question?

OpenStudy (aum):

Hang on. I found my mistake: (x + yi) + (2(x + yi) + 3i) = 9 x + yi + 2x + 2yi + 3i = 9 x + 2x + (yi + 2yi + 3i) = 9 3x + (3yi + 3i) = 9 3x + (y + 1)*3i = 9 (mistake in this line earlier). 3x = 9 and y + 1 = 0 x = 3 and y = -1 x + yi = 3 - i

OpenStudy (anonymous):

Thank you!